OpenStudy (anonymous):

Power (P) = Work (W) ÷ time (t) Joachim's body supplies a force of 450 N to run up a 20 meter hill in 15 seconds. How much power is involved in Joachim's run up the hill? A. 30 Watts B. 22.5 Watts C. 300 Watts D. 600 Watts

OpenStudy (anonymous):

d

OpenStudy (anonymous):

work =force*disctance=450*20=9000 time =15sec power=9000/15=600

OpenStudy (anonymous):

thanks i figured it out but i wanted to make sure thank you ^_^

OpenStudy (anonymous):

mention not :)